Write 1 page on magnetically levitated trains(maglev)
1
Expert's answer
2020-02-26T10:08:35-0500

Maglev (derived from magnetic levitation) is a system of train transportation that uses two sets of magnets, one set to repel and push the train up off the track, and another set to move the elevated train ahead, taking advantage of the lack of friction. Along certain "medium-range" routes (usually 320 to 640 km [200 to 400 mi]), maglev can compete favorably with high-speed rail and airplanes.[1]

The first patents that can be attributed to the devices of movement and suspension of cars on a magnetic cushion belong to the years 1905-1907 of the last century.

Transrapid 05

was the first maglev train with longstator propulsion licensed for passenger transportation. In 1979, a 908 m (2,979 ft) track was opened in Hamburg for the first International Transportation Exhibition (IVA 79).

The system of magnets that hold the car above the surface of the rails should ensure the stability of its suspension during rapid movement on high-speed sections. This is achieved not only by additional holding magnets, but also by the construction of paths. The Maglev rail rises above the ground for several purposes:

1. it should ensure that there are no accidental obstacles in the way of the car,

2. it should be very smooth with small curves,

3. this design ensures the independence of the route from the infrastructure of highways and urban structures.

Let's focus on point 2. We estimate the acceleration that occurs in the car when the track bends at a radius of R=5 km, at a speed of car V=600 km / h. The centrifugal acceleration is calculated using the formula "a=\frac{V^2}{R}". We determine "V=\frac {600 000 m}{3600 s}=167ms^{-1}" and "a=\frac {167^2}{5000m}=5.6ms^{-2}" which is more than half the body weight. The path should also be very flat in the vertical plane. If there are small but constantly occurring pits or ledges on the way, passengers will shake like on a bumpy road.

Various ideas were used for the stable position of the car in the suspension. They can be divided into several categories:

  1. electromagnetic suspension (EMS) systems;
  2. electrodynamic suspension (EDS) systems;
  3. permanent Magnet Passive Suspension.

Each of these designs has its own advantages and disadvantages.

Recently, high hopes are associated with the use of high-temperature superconductors with magnetic flux pinning [2].


The fact is that superconductors of the second kind have the ability to "freeze" the magnetic flux passing through their surface. If you try to remove them from the magnetic field in this state, a vortex current will flow over their surface (without resistance or dissipation), pulling them back. On the contrary, when approaching a magnet and increasing the magnetic field, they will tend to push away from it. The flow is freeze in a superconductor in the literal sense. A superconductor at a temperature above the critical (when is a ordinary conductor) is installed in the magnetic field of the rail at the required distance from it and cooled to a temperature below the critical temperature. After the transition to the superconducting state, the structure begins to work as indicated above.


[1] https://en.wikipedia.org/wiki/Maglev

[2] Hyung-Suk Han, Dong-Sung Kim,"Magnetic Levitation. Maglev technology and Applications", Springer, New York, 2016, p.56.